# RGB-Event based Pedestrian Attribute Recognition: A Benchmark Dataset and An Asymmetric RWKV Fusion Framework

### v1

@@ 1 Introduction
In this work, we first propose to recognize human attributes using a DVS346 event camera that outputs aligned frame-event streams. This is because event cameras have inherent advantages in high dynamic range, high temporal resolution, and low power consumption. However, event streams contain some noise and are less sensitive to static objects. Therefore, combining them with RGB frame cameras is necessary to achieve more robust and accurate results. Due to the lack of datasets in this area, we first propose a large-scale multi-modal RGB-Event pedestrian attribute recognition benchmark dataset, termed EventPAR. This dataset covers common human appearance attributes as well as six human emotional attributes, totaling 50 attributes and 100K paired RGB-Event samples, as shown in Fig. 1. These videos were collected under various challenging scenarios, including different lighting conditions, motion blur, object occlusion, and adverse weather conditions. In addition to natural challenging factors, we also artificially introduced some adversarial attack noises and degradation strategies, such as various types of noise and occlusion. To provide a solid benchmark for future research, we re-train and report 18 mainstream PAR algorithms, as shown in Table 3.
@@
1). We propose to recognize the pedestrian attributes using an event camera and construct a large-scale benchmark dataset termed EventPAR. It contains 100K aligned RGB frame and event streams, collected from multiple scenarios, illuminations, etc.
2). We propose a novel RWKV-based framework for the frame-event fusion based attribute recognition. A novel asymmetric RWKV fusion module fuses the RGB frame and event streams effectively and efficiently.
3). We build a large-scale benchmark for the EventPAR dataset by re-training 18 representative and strong PAR algorithms. This benchmark will be a good platform to boost the development of event-based attribute recognition further.
In addition, we also conduct extensive experiments on three PAR benchmark datasets, including the newly proposed EventPAR, MARS-Attribute and DukeMTMC-VID-Attribute dataset. More experiments fully validated the effectiveness of our proposed model for the frame-event pedestrian attribute recognition.
@@ 3.1 Protocols
In this work, we follow the guidelines below when constructing the EventPAR benchmark dataset: 1). Multi-Modal: Unlike existing frame-based PAR datasets, we utilize the DVS346 camera for data collection, obtaining spatiotemporal aligned multi-modal data comprising visible light and event streams. It is the first multi-modal PAR database, laying a solid data foundation for subsequent research. 2). Emotion Attributes: Unlike conventional PAR datasets that primarily focus on pedestrians’ appearance attributes, our new dataset incorporates six fundamental human emotional attributes, i.e., Happiness, Sadness, Anger, Surprise, Fear, and Disgust. This addition effectively fills the gap in the attribute recognition domain by addressing the lack of discussion on the emotional analysis dimension. 3). Diverse Scenes and Cross-Seasonal: Different from existing PAR datasets that involve short-term data collection, our data acquisition spanned several months, encompassing data from different seasons (summer and winter). Additionally, it covers a variety of scenes and weather conditions, such as daytime and nighttime scenes, as well as weather conditions like sunny and rainy days. These elements significantly enhance the diversity and challenge of the dataset, providing better validation of a model’s fitting ability and generalization capability. 4). Large Scale: It contains 100K pedestrian samples, 12 attribute groups, and 50 fine-grained pedestrian attributes. It is one of the largest PAR datasets, comparable to PA100K in scale. 5). Simulated Complex Real-world Environments: Our dataset is carefully designed to incorporate a wide range of challenges, such as variations in illumination, motion blur, object occlusion, and adverse weather conditions, all of which simulate the complex real-world difficulties encountered in pedestrian attribute recognition. To further increase the complexity and practical relevance of the dataset, we employ adversarial attack techniques [2, 32, 59] during data processing, adding an extra layer of difficulty. This approach not only enhances the robustness of pedestrian attribute recognition technologies but also ensures their applicability in real-world scenarios, as shown in Fig. 2.
Table 1: Comparison with existing PAR benchmark datasets. ##\##Att. is short for the number of Attributes. Deg and Adv is short for Degradation and Adversary, respectively.
Dataset Year ##\##Att. ##\##Samples Video Frame Event Emotion Deg Adv
@@
#05 Head Long Hair, Black Hair, Hat, Glasses, Mask, Scarf, Headphones
#06 Upper Body Short Sleeves, Long Sleeves, Shirt, Coat, Cotton-padded Coat
#07 Lower Body Trousers, Shorts, Jeans, Long Skirt, Short Skirt, Casual Pants, Dress
#08 Shoes Casual Shoes, Other Shoes
#09 Accessory Backpack, Shoulder Bag, Hand Bag, Umbrella, Cup, Cellphone
#10 Posture Walking, Running, Standing, Sitting
#11 Activity Smoking, Reading
#12 Emotion Happiness, Sadness, Anger, Surprise, Fear, Disgust
3.2 Data Collection and Annotation
Based on the aforementioned protocols, we collect the EventPAR dataset using a DVS346 event camera, which outputs both RGB frames and event streams. The entire data collection process spanned several months, covering multiple seasons from summer to winter. Approximately 20 students participated in the data collection and subsequent annotation process. As shown in Table 2, our dataset contains 100K spatio-temporal aligned RGB frames and event streams. Each sample in our dataset is annotated with 50 attributes and can be divided into 12 attribute groups, including Gender, Age, Body Size, Viewpoint, Head, Upper Body, Lower Body, Shoes, Accessory, Posture, Activity, and Emotion. More details can be found in Table 2.
By following the MSP60K dataset [20], we also introduced the data quality degradation strategy to evaluate the robustness of models against challenging factors, including changes in lighting, random occlusions, blurring, noise, and adversarial disturbance. Our dataset also exhibits a long-tail distribution, as depicted in Fig. 3 (a), which aligns with the attribute distribution typically observed in real-world data. It is evident that attributes such as Adult, Back Hair, and Normal Size are among the top-3 most frequent, while attributes like Smoking, Child, and Elderly rank at the bottom. We also provide the co-occurrence matrix of our EventPAR dataset as shown in Fig. 3 (b).
@@ 3.3 Benchmark Baselines
Based on our newly proposed dataset, we retrained and tested 18 mainstream PAR models, providing a solid baseline for future work on the EventPAR dataset. Specifically, the following algorithms were included: 1). CNN-based: DeepMAR [22],Zhou et al. [62], RethinkPAR [15], SSCNet [14], SSPNet [36]. 2). Transformer-based: DFDT [60], PARFormer [9]. 3). Mamba-based: MambaPAR [47], MaHDFT [46]. 4). Human-Centric Pre-Training Models for PAR: PLIP [65], HAP [57]. 5). Visual-Language Models for PAR: VTB [4], Label2Label [25], PromptPAR [45], SequencePAR [19].
@@ 4.1 Overview
To enhance recognition accuracy and robustness, in this paper, we propose a video-based multi-modal pedestrian attribute recognition framework, as illustrated in Fig. 4. Our method leverages event data as the auxiliary modality to compensate for information loss in the RGB modality under challenging environments, thereby improving the stability of recognition. With the powerful modeling capabilities of Vision-RWKV, we propose an OTN-RWKV guided multimodal fusion framework. Concretely, we first employ the Visible-RWKV encoder and Event-RWKV encoder to extract the specific features from RGB and event data, respectively. The extracted features are subsequently processed by the proposed OTN-RWKV fusion module to facilitate the interaction and integration of multimodal information. The resulting fused representation is then passed to a classifier for accurate attribute prediction. Please refer to the following for more detailed information.
@@ 4.2 Network Architecture
$\bullet$ Input Encoding : In this work, event modality is used to assist the RGB video in overcoming information loss in complex environments. Given a sequence of an RGB video denote as $X_{r}\in\mathbb{R}^{T\times H\times W\times C}$, where $T$ is the total number of frames, and $H$, $W$, $C$ correspond to the height, width, and the number of channels of each frame, respectively. We take event streams as =$\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$, each event point $e_{j}$ can be denoted as $\{x,y,t,p\}$, representing spatial coordinates, timestamps, and polarity, respectively. To align with the RGB sequence, we stack the event streams into event frames that are temporally synchronized with the RGB frames based on exposure times. Therefore, we can obtain the event frame sequence $X_{e}\in\mathbb{R}^{T\times H\times W\times C}$.
@@
$\bullet$ RWKV Encoder: In this paper, we employ Vision-RWKV [8] as our visual encoder, which adopts a block-stacked design with $L$ layers to enhance its encoding capability, where each block consists of a Spatial-Mix module and a Channel-Mix module. In each layer, RGB tokens $X_{r}\in\mathbb{R}^{T\times N\times C}$ and event visual tokens $X_{e}\in\mathbb{R}^{T\times N\times C}$ are first fed into the Spatial-Mix module, which performs linear complexity global attention computation. Specifically, as shown in the top right of Fig. 4, the input visual tokens first undergo a token Shift operation to capture local context information before being fed into three parallel linear layers to obtain the matrices $R_{s}$,$K_{s}$,$V_{s}\in\mathbb{R}^{T*N\times C}$. The formulas are as follows,
@@
where the Q-Shift function is a specialized token shift mechanism specifically designed for vision tasks. After that, $K_{s}$ and $V_{s}$ are used to calculate the global attention result $wkv\in\mathbb{R}^{T*N\times C}$ by the Bi-WKV module, which is a linear complexity bidirectional attention mechanism. Subsequently, the $wkv$ is multiplied with $\sigma(R)$, which controls the output $O_{i}$ probability,
$O_{i}$ $=(\sigma(R_{s})\odot whv)W_{O},$ (2)
[…]
@@
[…]
the gating mechanism $\sigma(R_{c})$ is used to modulate the output $O^{\prime}_{i}$. To maintain information integrity and promote gradient flow, residual connections are introduced to fuse the module outputs, thereby effectively alleviating the vanishing gradient problem. This process can be expressed as,
@@
$\bullet$ OTN-RWKV: Due to the asynchronous nature of event cameras, the event streams tend to be highly dense, resulting in stacked event frames containing a significant amount of redundant information. This redundancy not only increases computational overhead but also causes information ambiguity, thereby diminishing the effectiveness of event data as auxiliary information. To address this issue, we propose the OTN-RWKV fusion module, which first applies the similarity-based filtering on the visual tokens output $O_{e}^{\prime}$ of event modality, removing highly redundant tokens and achieving one-to-N multimodal fusion. Specifically, a similarity matrix is employed to identify the top-$K$ most similar token pairs, from which only the most representative tokens are retained. This approach enhances data efficiency and improves the quality of multimodal fusion. The implementation process is as follows,
@@
where $O_{e}^{{}^{\prime\prime}}\in\mathbb{R}^{N^{\prime}\times C}$ represents the tokens for the informative filtered event. After filtering, both modalities are aligned to have the same number of tokens, facilitating effective multimodal fusion.
Although traditional fusion methods (e.g., concatenation, addition, and 1×1 convolution) generally exhibit strong compatibility and generalization across different models, they frequently result in information loss, limiting their effectiveness in preserving fine-grained multimodal details. Inspired by the concept of cross attention and LCR[54], in this work, we further propose a novel interactive fusion strategy. As shown in Fig. 5, we first apply the Q-shift operation separately to the two types of visual tokens, corresponding to Eq.1 and Eq. 3, respectively. Subsequently, $V_{s}$, $R_{c}$, and $K_{c}$ are used to compute the cross-attention results through a linear complexity bidirectional attention mechanism,
@@ 4.3 Loss Function
$\mathcal{L}_{wce}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\omega_{j}\left(y_{ij}\log(p_{ij})+(1-y_{ij})\log(1-p_{ij})\right),$ (8)
@@ 5.1 Datasets & Evaluation Metric
In this study, we conduct a comprehensive benchmark of 18 pedestrian attribute recognition methods, representing the most important models in the field of pedestrian attribute recognition. Additionally, the performance of our method is compared with existing state-of-the-art (SOTA) pedestrian attribute recognition (PAR) methods across our benchmark and two publicly available datasets: MARS-Attribute Dataset [61] and DukeMTMC-VID-Attribute Dataset [34].
@@
Table 3: Comparison with public methods on our datasets. The first and second are shown in red and blue, Size indicates the size of checkpoint file, and ##\##MS means the GPU memory used for optimization, Test time refers to the time spent on the test set.
@@
Methods Publish Code mA Acc Prec Recall F1 Test Time Params (MB) Flops (GB) Size (MB) ##\##MS (MB)
#01  DeepMAR [22] ACPR15 URL 71.15 72.39 86.42 78.93 82.51 46s 23 4 181 846
#02  ALM  [39] ICCV19 URL 71.31 68.80 78.82 75.97 77.37 21s 17 45 66 484
#03  Strong Baseline [15] arXiv20 URL 73.75 70.63 81.34 78.26 79.32 91s 24 4 91 788
#04  RethinkingPAR [15] arXiv20 URL 71.23 69.00 80.80 77.16 78.94 57s 24 4 91 105
#05  SSCNet [14] ICCV21 URL 72.12 69.95 78.36 80.10 79.22 24s 24 3 90 854
#06  VTB [4] TCSVT22 URL 76.96 73.81 80.74 83.15 81.67 441s 93 15 333 114
#07  Label2Label [25] ECCV22 URL 72.47 69.54 78.75 79.14 78.55 43s 66 12 829 3788
#08  DFDT [60] EAAI22 URL 70.14 70.95 78.71 81.82 80.24 193s 88 26 669 1884
#09  Zhou et al. [62] IJCAI23 URL 70.16 67.41 79.76 76.40 78.04 76s 140 15 536 4176
#10  PARFormer [9] TCSVT23 URL 75.33 74.21 79.33 85.82 82.09 732s 195 68 756 11634
#11  SequencePAR [19] arXiv23 URL 73.96 73.40 81.47 81.17 81.11 275s 466 413 1776 11263
#12  VTB-PLIP  [65] arXiv23 URL 73.72 71.50 79.97 80.57 79.99 26s 31 8 591 1614
#13  Rethink-PLIP [65] arXiv23 URL 57.91 62.78 78.27 70.81 74.35 44s 24 4 144 80
#14  PromptPAR [45] TCSVT24 URL 74.92 71.79 78.33 83.01 80.31 977s 8 3 1200 4435
#15  SSPNet  [63] PR24 URL 71.71 69.41 80.46 77.46 78.93 32s 26 7 100 1240
#16  HAP [57] NIPS24 URL 77.33 74.13 83.99 80.56 82.24 92s 86 17 327 2683
#17  MambaPAR [47] arXiv24 URL 72.28 70.57 77.99 81.70 79.50 18s 26 1 98 116
#18  MaHDFT [46] arXiv24 URL 75.54 72.27 80.28 82.19 80.59 158s 158 3978 431 187
              Ours - - 79.32 76.00 82.37 84.55 83.22 68s 107 39 429 2200
              Ours+E - - 87.66 84.78 89.03 89.38 89.07 361s 201 77 836 2312
5.2 Implementation Details
5.3 Comparison on Public Datasets
$\bullet$ Results on EventPAR Dataset.  We collect and analyze public PAR methods from 2015 to 2024 on the EventPAR dataset as shown in Table 3, methods such as HAP [57], SequencePAR [19], and PARFormer [9] have demonstrated superior performance. Among them, the HAP method achieved outstanding results, with mA, Acc, Prec, Recall, and F1 scores of 77.33, 74.13, 83.99, 80.56, and 82.24, respectively. Our method achieves strong performance with mA, Acc, Prec, Recall, and F1 scores of 79.32, 76.00, 82.37, 84.55, and 83.22, respectively, outperforming existing approaches across most evaluation metrics. Furthermore, by incorporating the event auxiliary modality, our model further improves these scores to 87.66, 84.78, 89.03, 89.38, and 89.07, achieving the best overall results. These experimental findings on the EventPAR dataset not only confirm the effectiveness of our method in pedestrian attribute recognition tasks but also highlight the significant advantage of integrating event modality in enhancing model robustness under complex real-world conditions.
$\bullet$ Results on MARS-Attribute Dataset [61] As shown in Table 4, to further evaluate the generalization and effectiveness of our method, we conducted comprehensive experiments on another public dataset, comparing it with several state-of-the-art approaches. While our method did not consistently achieve the top performance across all metrics, it demonstrated strong competitiveness, achieving 73.74, 80.43, 84.63, and 82.18 in Accuracy, Precision, Recall, and F1, respectively. These results reflect a balanced performance with high accuracy and effective attribute capture.
$\bullet$ Results on DukeMTMC-VID-Attribute Dataset [34]. Our method achieves performance comparable to mainstream approaches on this dataset, with mA, Accuracy, Precision, Recall, and F1 of 68.19, 82.83, 77.30, 79.38, and 78.32, respectively. Although the overall performance is slightly lower than that of the VTB method, our approach still demonstrates solid generalization ability and stability. Notably, it achieves competitive results while maintaining relatively low computational complexity, validating the effectiveness and feasibility of our proposed method. These findings suggest that our method remains competitive in certain real-world scenarios and has the potential for further improvement through optimization of the model architecture and fusion strategy.
@@ 5.3 Comparison on Public Datasets
Accuracy Precision Recall F1 Accuracy Precision Recall F1
@@
Ours RWKV-B 73.74 80.43 84.63 82.18 68.19 82.83 77.30 79.38
Table 5: Comparison of different input settings and filtering threshold is set to 0.95.

### v2

@@ 1 Introduction
In this work, we first propose to recognize human attributes using a DVS346 event camera that outputs aligned frame-event streams. This is because event cameras have inherent advantages in high dynamic range, high temporal resolution, and low power consumption. However, event streams contain some noise and are less sensitive to static objects. Therefore, combining them with RGB frame cameras is necessary to achieve more robust and accurate results. Due to the lack of datasets in this area, we first propose a large-scale multi-modal RGB-Event pedestrian attribute recognition benchmark dataset, termed EventPAR. This dataset covers common human appearance attributes as well as six human emotional attributes, totaling 50 attributes and 100K paired RGB-Event samples, as shown in Fig. 1. These videos were collected under various challenging scenarios, including different lighting conditions, motion blur, object occlusion, and adverse weather conditions. In addition to natural challenging factors, we also artificially introduced some adversarial attack noises and degradation strategies, such as various types of noise and occlusion. To provide a solid benchmark for future research, we re-train and report 17 mainstream PAR algorithms, as shown in Table 3.
@@
1). We propose to recognize the pedestrian attributes using an event camera and construct a large-scale benchmark dataset termed EventPAR. It contains 100K aligned RGB frames and event streams, collected from multiple scenarios, illuminations, etc.
2). We propose a novel RWKV-based framework for frame-event fusion based attribute recognition. A novel asymmetric RWKV fusion module fuses the RGB frame and event streams effectively and efficiently.
3). We build a large-scale benchmark for the EventPAR dataset by re-training 17 representative and strong PAR algorithms. This benchmark will be a good platform to boost the development of event-based attribute recognition further.
In addition, we also conduct extensive experiments on three PAR benchmark datasets, including the newly proposed EventPAR, MARS-Attribute, and DukeMTMC-VID-Attribute datasets. More experiments fully validated the effectiveness of our proposed model for the frame-event pedestrian attribute recognition.
@@ 3.1 Protocols
In this work, we follow the guidelines below when constructing the EventPAR benchmark dataset: 1). Multi-Modal: As shown in Table 2, unlike existing frame-based PAR datasets, we utilize the DVS346 camera for data collection, obtaining spatiotemporal aligned multi-modal data comprising visible light and event streams. It is the first multi-modal PAR database, laying a solid data foundation for subsequent research. 2). Emotion Attributes: Unlike conventional PAR datasets that primarily focus on pedestrians’ appearance attributes, our new dataset incorporates six fundamental human emotional attributes, i.e., Happiness, Sadness, Anger, Surprise, Fear, and Disgust. This addition effectively fills the gap in the attribute recognition domain by addressing the lack of discussion on the emotional analysis dimension. The static and dynamic presentations are shown on the right side of Fig. 2. 3). Diverse Scenes and Cross-Seasonal: Different from existing PAR datasets that involve short-term data collection, our data acquisition spanned several months, encompassing data from different seasons (summer and winter). Additionally, it covers a variety of scenes and weather conditions, such as daytime and nighttime scenes, as well as weather conditions like sunny and rainy days, as shown in Fig. 2 (a, b). These elements significantly enhance the diversity and challenge of the dataset, providing better validation of a model’s fitting ability and generalization capability. 4). Large Scale: It contains 100K spatio-temporally aligned pedestrian samples of RGB frames and event streams, 12 attribute groups, and 50 fine-grained pedestrian attributes. It is one of the largest PAR datasets, comparable to PA100K in scale, as shown in Table 2. 5). Simulated Complex Real-world Environments: Our dataset is carefully designed to incorporate a wide range of challenges, such as variations in illumination, motion blur, object occlusion, and adverse weather conditions, all of which simulate the complex real-world difficulties encountered in pedestrian attribute recognition. To further increase the complexity and practical relevance of the dataset, we employ adversarial attack techniques [2] during data processing, adding an extra layer of difficulty, as shown in Fig. 2 (c).
Moreover, EventPAR exhibits a real-world long-tail distribution, as shown in Fig. 3 (a), where frequent attributes include Adult, Back Hair, and Normal Size while rare ones such as Smoking, Child, and Elderly appear less often. The attribute co-occurrence matrix is shown in Fig. 3 (b).
Table 1: Comparison with existing PAR benchmark datasets. #\#Att. is short for the number of Attributes. Deg and Adv is short for Degradation and Adversary, respectively.
Dataset Year #\#Att. #\#Samples Video Frame Event Emotion Deg Adv
@@
#05 Head
Long Hair, Black Hair, Hat, Glasses, Mask, Scarf, Headphones
#06 Upper Body
Short Sleeves, Long Sleeves, Shirt, Coat, Cotton-padded Coat
#07 Lower Body
Trousers, Shorts, Jeans, Long Skirt, Short Skirt, Casual Pants, Dress
#08 Shoes
Casual Shoes, Other Shoes
#09 Accessory
Backpack, Shoulder Bag, Hand Bag, Umbrella, Cup, Cellphone
#10 Posture
Walking, Running, Standing, Sitting
#11 Activity
Smoking, Reading
#12 Emotion
Happiness, Sadness, Anger, Surprise, Fear, Disgust
@@ 3.2 Benchmark Baselines
Based on our newly proposed dataset, we retrained and tested 17 mainstream PAR models, providing a solid baseline for future work on the EventPAR dataset. Specifically, the following algorithms were included: 1). CNN-based: DeepMAR [22],Zhou et al. [59], RethinkPAR [15], SSCNet [14], SSPNet [35]. 2). Transformer-based: DFDT [57], PARFormer [9]. 3). Mamba-based: MambaPAR [46], MaHDFT [45]. 4). Human-Centric Pre-Training Models for PAR: PLIP [62]. 5). Visual-Language Models for PAR: VTB [4], Label2Label [25], PromptPAR [44], SequencePAR [19].
@@ 4.1 Overview
To improve recognition accuracy and robustness, this paper proposes a video-based multimodal pedestrian attribute recognition framework shown in Fig. 4. Event data is used to supplement RGB information loss, enhancing stability. Features are extracted via Visible-RWKV and Event-RWKV encoders, fused through the OTN-RWKV module, and then classified for attribute prediction. Please refer to the following for more detailed information.
@@ 4.2 Network Architecture
$\bullet$ Input Encoding: In this work, event modality is used to assist the RGB video in overcoming information loss in complex environments. Given a sequence of an RGB video denote as $X_{r}\in\mathbb{R}^{T\times H\times W\times C}$, where $T$ is the total number of frames, and $H$, $W$, $C$ correspond to the height, width, and the number of channels of each frame, respectively. We take event streams as $\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$, each event point $e_{j}$ can be denoted as $\{x,y,t,p\}$, representing spatial coordinates, timestamps, and polarity, respectively. To align with the RGB sequence, we stack the event streams into event frames that are temporally synchronized with the RGB frames based on exposure times. Therefore, we can obtain the event frame sequence $X_{e}\in\mathbb{R}^{T\times H\times W\times C}$.
@@
$\bullet$ RWKV Encoder: In this paper, we employ Vision-RWKV [8] as our visual encoder, which adopts a block-stacked design with $L$ layers to enhance its encoding capability, where each block consists of a Spatial-Mix module and a Channel-Mix module. In each layer, RGB tokens $X_{r}\in\mathbb{R}^{T\times N\times C}$ and event visual tokens $X_{e}\in\mathbb{R}^{T\times N\times C}$ are first fed into the Spatial-Mix module, which performs linear complexity global attention computation. Specifically, as shown in the top right of Fig. 4, the input visual tokens first undergo a token Shift operation to capture local context information before being fed into three parallel linear layers to obtain the matrices $R_{s}$$K_{s}$$V_{s}\in\mathbb{R}^{T*N\times C}$. The formulas are as follows:
@@
where the Q-Shift function is a specialized token shift mechanism specifically designed for vision tasks. After that, $K_{s}$ and $V_{s}$ are used to calculate the global attention result $wkv\in\mathbb{R}^{T*N\times C}$ by the Bi-WKV module, which is a linear complexity bidirectional attention mechanism. Subsequently, the wkvwkv is multiplied with $\sigma(R)$, which controls the output $O_{i}$ probability,
$O_{i}$ $=(\sigma(R_{s})\odot wkv)W_{O},$ (2)
$wkv$ $=\text{Bi-WKV}(K_{s},V_{s}),\quad i\in(r,e),$
@@
$R_{c}$ $=\text{Q-Shift}_{R}(O_{i})W_{R},$ (3)
$K_{c}$ $=\text{Q-Shift}_{K}(O_{i})W_{K},$
the gating mechanism $\sigma(R_{c})$ is used to modulate the output $O^{\prime}_{i}$. To maintain information integrity and promote gradient flow, residual connections are introduced to fuse the module outputs, thereby effectively alleviating the vanishing gradient problem. This process can be expressed as:
@@
$\bullet$ OTN-RWKV: Due to the asynchronous nature of event cameras, event frames contain a large amount of redundant information, increasing computational burden and reducing the effectiveness of auxiliary data. To address this, we propose the OTN-RWKV fusion module, which uses a similarity matrix to identify the top-$K$ most similar token pairs from the event tokens output $O_{e}^{\prime}$, retaining only the most representative tokens. This method improves computational efficiency and enhances the quality of multimodal fusion. The implementation process is as follows:
@@
where $O_{e}^{{}^{\prime\prime}}\in\mathbb{R}^{N^{\prime}\times C}$ represents the tokens for the informative filtered event. After filtering, the number of tokens in both modalities is aligned to facilitate fusion.
Although traditional fusion methods (e.g., concatenation, addition, and 1×1 convolution) offer strong compatibility, they often cause information loss, limiting the preservation of fine-grained multimodal details. Inspired by the concept of cross attention and LCR[53], in this work, we further propose a novel interactive fusion strategy. As shown in Fig. 5, we first apply the Q-shift operation separately to the two types of visual tokens, corresponding to Eq.1 and Eq. 3, respectively. Subsequently, $V_{s}$, $R_{c}$, and $K_{c}$ are used to compute the cross-attention results through a linear complexity bidirectional attention mechanism,
@@ 4.3 Loss Function
$\small\mathcal{L}_{wce}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\omega_{j}\left(y_{ij}\log(p_{ij})+(1-y_{ij})\log(1-p_{ij})\right),$ (8)
@@ 5.1 Datasets & Evaluation Metric
This study benchmarks 17 pedestrian attribute recognition methods and compares our method with state-of-the-art approaches on the benchmark and public datasets MARS-Attribute [58] and DukeMTMC-VID-Attribute [33].
@@
Table 3: Comparison with public methods on our datasets. Size indicates the size of checkpoint file, and #\#MS means the GPU memory used for optimization, Test time refers to inference time.
@@
Methods Publish mA Acc Prec Recall F1 Test Time Params (MB) Flops (GB) Size (MB) #\#MS (MB)
#01  DeepMAR [22] ACPR15 66.57 69.53 74.90 88.54 81.57 64s 23 25 181 592
#02  ALM  [38] ICCV19 57.18 64.17 75.59 73.20 74.38 813s 17 8 66 785
#03  Strong Baseline [15] arXiv20 73.75 61.86 67.23 80.78 75.43 47s 24 25 91 576
#04  RethinkingPAR [15] arXiv20 81.37 80.84 86.31 87.57 86.93 44s 24 24 91 580
#05  SSCNet [14] ICCV21 63.10 66.07 72.72 83.22 77.62 49s 24 9 90 588
#06  VTB [4] TCSVT22 88.41 83.83 87.89 89.31 88.53 243s 93 66 333 525
#07  Label2Label [25] ECCV22 72.49 74.01 86.60 79.02 82.19 398s 66 37 829 650
#08  DFDT [57] EAAI22 61.71 63.14 79.17 70.63 74.66 287s 88 37 669 2584
#09  Zhou et al. [59] IJCAI23 56.46 60.89 73.37 73.62 73.50 42s 234 24 536 701
#10  PARFormer [9] TCSVT23 83.12 80.48 85.14 88.41 86.53 438s 195 205 756 2481
#11  SequencePAR [19] arXiv23 86.27 84.42 88.81 89.12 88.83 1497s 466 577 1776 9901
#12  VTB-PLIP  [62] arXiv23 67.25 68.37 77.75 79.72 78.37 229s 31 22 591 501
#13  Rethink-PLIP [62] arXiv23 68.75 70.03 81.82 78.04 79.89 37s 21 23 144 481
#14  PromptPAR [44] TCSVT24 86.51 82.27 86.35 89.36 87.64 1312s 8 5 1200 2296
#15  SSPNet  [60] PR24 66.92 67.49 78.73 76.90 77.80 69s 40 26 100 1898
#16  MambaPAR [46] arXiv24 50.01 42.32 54.81 57.31 55.63 105s 0.64 0.41 98 802
#17  MaHDFT [45] arXiv24 50.43 44.98 59.10 59.70 58.57 292s 164 6033 499 1313
             OTN-RWKV (Ours, RGB Only) - 79.32 76.00 82.37 84.55 83.22 68s 107 39 429 2200
             OTN-RWKV (Ours, RGB+Event) - 87.70 84.94 89.15 89.48 89.18 361s 201 77 836 2312
@@
Acc Prec Recall F1 Acc Pren Recall F1
@@
OTN-RWKV(Ours) RWKV-B 73.21 85.63 81.53 83.22 73.15 84.45 82.16 82.78
5.2 Implementation Details
5.3 Comparison on Public Datasets
$\bullet$ Results on EventPAR Datasets.  We collect and analyze public PAR methods from 2015 to 2025 on the EventPAR dataset as shown in Table 3, methods such as SequencePAR [19], and PARFormer [9] have demonstrated superior performance. Our method achieves good results using only RGB data, with mA, Acc, and F1 scores of 79.32, 76.00, and 83.22 respectively. Adding the event modality further improves these to 87.66, 84.78, and 89.07, achieving the best overall performance and demonstrating the significant benefit of event data for model robustness in complex environments. Because PAR methods from 2025 have not released their code, experiments cannot be conducted.
$\bullet$ Results on MARS-Attribute Datasets [58] As shown in Table 4, comparing our method with several state-of-the-art approaches. Our method achieves the best performance on most evaluation metrics, achieving 73.21, 85.63, 81.53, and 83.22 in Accuracy, Precision, Recall, and F1, respectively. These results reflect a balanced performance with high accuracy and effective attribute capture.
$\bullet$ Results on DukeMTMC-VID-Attribute Datasets [33]. As shown in Table 4, we also compared our method with several mainstream approaches on this dataset, achieving strong results with accuracy, precision, recall, and F1 scores of 73.15, 84.45, 82.16, and 82.79, respectively. Our method shows stable performance across diverse data environments, confirming its effectiveness and adaptability, with strong application potential.
5.4 Ablation Study
In this section, we conduct a comprehensive ablation study on the core modules of EventPAR, including different input configurations, various backbone architectures, fusion strategies and event data aggregation methods.
$\bullet$ Analysis of different input settings. In this study, We varied RGB input counts and ran extensive experiments as shown in Table 5. RGB or Event alone performed well, with RGB slightly worse due to noise. Combining both improved results, showing complementarity. Too many RGB inputs hurt performance due to redundancy and noise, burdening feature extraction and weakening semantic focus.
Table 5: Comparison of different input settings.